OpenStudy (anonymous):

Bob has some 10 lb weights and some 3 lb weights. Together, all his weights add up to 50 lbs. If x represents the number of 3 lb weights and y represents the number of 10 lb weights, which equation can be used to find the number of each type of weight Bob has?

OpenStudy (boldjon):

50 + 10y = 3y

OpenStudy (tgannon998):

10x +3y=50

OpenStudy (medicaldoctor):

3x + 10y = 50

OpenStudy (medicaldoctor):

@Tgannon998 and @boldjon I do not see how you came up with your answers from the information given?

OpenStudy (tgannon998):

same way you came up with yours.

OpenStudy (medicaldoctor):

Right but you have to be careful because your answer is wrong. Your variables are mixed up.

OpenStudy (tgannon998):

it doesnt matter what is labeled X or Y.

OpenStudy (medicaldoctor):

Actually it does matter because the problem specifically says what to label them. y = 10 lbs and x = 3 lbs

OpenStudy (tgannon998):

It does the same thing.

OpenStudy (medicaldoctor):

@mathmale am I making sense here? And it doesn't actually because the problem would be different if you had what it says and the variables you put down. Just be careful.

OpenStudy (medicaldoctor):

Plugging in information correctly is key to Algebra, and simple math.

OpenStudy (tgannon998):

you still get the same answer.

OpenStudy (medicaldoctor):

No you will not because you are plugging in 3 times the amount of 10 lbs and 10 times the amount of 3 lbs.

OpenStudy (medicaldoctor):

Since we have 3x for our weights the amount of weight total of the 3lb weights is going to be three times. This is why we need to use x.

OpenStudy (texaschic101):

it has to be 3x + 10y = 50....because x represents 3 lb....and y represents 10 lb

OpenStudy (medicaldoctor):

We have 10y because the amount of weight total will be ten times the amount of weights. This is why y is important. And exactly @texaschic101